OpenStudy (anonymous):

find two numbers whose sum is 46, and whose difference is 32

OpenStudy (anonymous):

why not guess?

OpenStudy (anonymous):

i guess 10 and 44 but that is not right because they add up to 54 not 46 try starting with a smaller number

OpenStudy (anonymous):

\[a + b = 46\] \[a-b= 32\]

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Rearrange the equations and solve for either A or B and plug it into the other

OpenStudy (anonymous):

a syste m of 2 equations :D

OpenStudy (anonymous):

why does everyone need algebra?

OpenStudy (anonymous):

just do it !!!

OpenStudy (anonymous):

algebra makes things easier right

OpenStudy (anonymous):

!!!!

OpenStudy (anonymous):

hell now just guess bet you get it on the second try

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

we have a sum of x and a difference of y. for any problem of this sort we take the difference and divide it by 2 (9) and subtract and add it to one half the sum (23) so we have the larger number as 23+9=32 and the smaller number as 23-9=14.... and just to check 32+ 14 = 46 . 32 - 14 = 18 Magical!!